OpenStudy (anonymous):

Given a cube with a volume of 27 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube? 3cm, 6cm, 9cm, or 12cm

OpenStudy (hoblos):

area of pyramid = (area of cube)/3 = 27/3 = 9cm3

OpenStudy (anonymous):

thank you

OpenStudy (hoblos):

welcome :)